The firm’s cost of equity capital is 18%, the market value of firm’s equity is $8 million, the firm’s cost of debt is 9%, and the market value of debt is $4 million. The firm is considering a new investment with an expected rate of return of 17%. This project is 30% riskier than the firm’s average operations. The risk-free rate of return is 5%, the variance of the market return is 0.08. Is the project profitable? 


1
Expert's answer
2022-01-25T06:47:34-0500

"WACC=18\times\frac{8}{12}+9\times\frac{4}{12}=12+3=15"


"CARM=5+\frac{0.3}{0.08}(15-5)=42.5"


17<42.5


the project is profitable